youre looking at an intervestment that will pay $1200 in 7 years if you invest $800 today. write out the formula you would use in determining the interest rate for this specific problem.
Answer #1

Future value =present value (1+i)^n

1200 = 800 (1+i)^7

1200/800 = (1+i)^7

    1.5 = (1+i)^7

multiplying both sides by ^1/7

(1.5)^1/7 = (1+ i)

    1.0596= 1+ i

    i = 1.0596 -1

        = .0596 or 5.96%